OpenStudy (anonymous):

Multiply or divide. ***PROBLEM REQUIRES EXPONENTS. SEE COMMENTS FOR PROBLEM.***

OpenStudy (anonymous):

\[(-x ^{3}) (-x)\]

OpenStudy (anonymous):

Multiply you get x^4

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Parenthesis is honestly another method of multiplying. Such as (4)(2) = 4*2. Through that logic it can be applied to variables.

OpenStudy (anonymous):

so the answer is \[x ^{4}\]

OpenStudy (anonymous):

The most basic way i can rewrite your problem would be (-1)*(x^3)*(-1)*(x)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Two negatives yield a positive solution

OpenStudy (anonymous):

are you positive??? because this is for a test retake and i need to make sure i do these right

OpenStudy (anonymous):

When multiplying by the same base you can add it to the exponent. 2*2*2*2 = 2^4 the same could be applied to your problem. And I'm sure the answer is x^4

OpenStudy (anonymous):

I don't know where u get the 2 or the 4 because there are no numbers in this problem. just x's.

OpenStudy (anonymous):

and the exponent 3

OpenStudy (anonymous):

I'm using the numbers as an example. x^3 = x*x*x

OpenStudy (anonymous):

Your problem is basically (-1)*(-1)*x*x*x*x = x^4